OpenStudy (anonymous):

[5+(x-y)] [5-(x-y)] Answer it please?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

just multiply it out; but watch the -(x-y) part; it distributes into (-x+y) 5+x-y 5-x+y -------- 25+5x -5y -5x -x^2 +xy +5y +5xy -y^2 ------------------------- 25 -x^2 +6xy -y^2

OpenStudy (anonymous):

use the formula (a+b)(a-b)=a^2-b^2 therefore a=5 b=x-y so 25-(x-y)^2